OpenStudy (anonymous):

1-2/3(4-5x)=5 So how do you get 3-2(4-5x)=15 out from that?? Help please?

OpenStudy (anonymous):

I dont fully understand what you are asking but im looking at it as x=2 i did the pemdas

OpenStudy (anonymous):

It's a proof that I have to figure out. I know the answer to the entire thing is x=2, but I have to figure out every step. I just don't understand how the Given is 1-2/3(4-5x)=5 and how they come up with the next step as 3-2(4-5x)=15

OpenStudy (anonymous):

3-2(4-5x)=15 Distribute the -2 3-8+10x=15 Combine like terms -5+10x=15 Add 5 to both sides 10x=20 Divide by 10 x=2 i just lookd it up...

OpenStudy (ghazi):

\[1-\frac{ 2 }{ 3(4-5x)}=15\] after taking LCM you'll have \[\frac{ 3(4-5x)-2 }{ 3(4-5x) }=15\] then use cross multiply....\[\frac{ 12-15x-2 }{ 12-15x }=15\] then \[\frac{ 10-15x }{ 1 }=\frac{ 45(4-5x) }{ 1 }\] then

OpenStudy (ghazi):

\[5*(2-3x)=45(4-5x).....(2-3x)= 9(4-5x)\]

OpenStudy (anonymous):

ghazi u r one smart pickle o.e

OpenStudy (ghazi):

:) well i was seeing you from a long time here....so thought to initiate :)

OpenStudy (ghazi):

\[2-3x= 36-45x\] then \[45x-3x=36-2=34......42x=34\] solve for x